LASER INDUCED FLUORESCENCE SPECTROSCOPY OF THE C_{4}H AND C_{4}D RADICALS

\underline{K. HOSHINA} , H. KOHGUCHI, AND Y. ENDO, Department of Pure and Applied Sciences, College of Arts and Sciences, The University of Tokyo, 153 Tokyo, Japan; Y.OHSHIMA, Department of Chemistry, Graduate School of Science, Kyoto University, 606-01 Kyoto, Japan.

Laser-induced-fluorescence spectra of the C_{4}H and C_{4}D radicals have been observed for the first time. The molecules were generated in a supersonic jet by an electric discharge of C_{2}H_{2} or C_{2}D_{2} diluted in Ar. Fourteen and thirteen vibronic bands were observed for C_{4}H and C_{4}D respectively, in the near UV region(24000 - 25000 cm^{−1}). The bands were assigned to the ^{2}Σ-^{2}Σ or ^{2}Π-^{2}Σ type, and several ^{2}Σ-^{2}Σ bands were subjected to rotational analyses. The determined rotational constants for the lower state agree with those for C_{4}H/C_{4}D in the ground vibrational level in the X^{2}Σ state, which were reported in a previous microwave study ^{a} ^{b}. The spin-orbit splittings in the ^{2}Π-^{2}Σ bands are 16 and 12 cm^{−1} for C_{4}H and C_{4}D, respectively, leading a conclusion that the excited electronic state for the observed band system is ^{2}Π. The present observation is consistent with a recent CASSCF study^{c}, which predicted the B^{2}Π state with the excitation energy of 24500 cm^{−1}.

aC.A. Gottlieb, E.W. Gottlieb, P. Thaddeus, and M. Kawamura, Astrophys. J. 275, 916 (1983).

bD.R. Woodward, J.C. Person, C.A Gottlieb, and C.A Thaddeus, Astrophys. J. 333, L29 (1988).

cA.L. Sobolewski and L. Adamowicz, J. Chem. Phys. 102, 394 (1995).